Calculate the average excess of fitness for each allele. Show your work.



Given the following allele frequency data, answer the following questions. 


Generation 1: p=0.6, q=0.4; population mean fitness =0.9 

Generation 2: p'=0.54, q'=0.46 


Calculate the average excess of fitness for each allele. Show your work.
Answer #1

Answer :)

To calculate the average excess of fitness (αA1), use the following formula:

Δp = p x (αA1 / wbar)

The symbol wbar (overbar of w) is the average fitness of the population.

Fitness of allele p = 0.6

Δp = 0.6 – 0.54

Δp = 0.06

0.06 = 0.6 x (αA1 /0.9)

(αA1 /0.9) = 0.1

αA1 = 0.09

Therefore, average excess of fitness for the allele p is 0.09.

For allele q:

Δq = q x (αA2 / wbar)

Δq = 0.46 – 0.4

Δq = 0.06

0.06 = 0.4 x (αA2 /0.9)

(αA2 /0.9) = 0.06/0.4

(αA2 /0.9) = 0.15

αA2 = 0.135

Therefore, average excess of fitness for the allele q is 0.135.

PLEASE RATE THE ANSWER